# Steering AI-Driven Personalization of Scientific Text for General Audiences

### v1

@@ 3.1.1. Translation Panel
The translation panel housed the main personalization features (Figure 1 A, B, C). At the top, the personalized title of the scientific article being viewed is displayed, with a box below presenting the personalized translation. A slider, labeled from weak to strong personalization, allowed participants to adjust the degree of personalization. The current personalization degree, a numeric value between 0 and 100, was shown just beneath the slider. Participants were instructed to adjust the slider to their preferred degree value and click Translate at the bottom. This action triggered a call to an LLM, which generated and displayed a personalized translation in the box above.
@@ 3.1.2. History Box
The history box on the right allowed participants to revisit, review, and compare translations they had seen during the session. Each new translation was automatically added to the stack, labeled with its personalization degree and the time of generation. Participants could click on any past translation to view it, which would update the translation panel accordingly. They could also mark a translation as a favorite by clicking on the star icon beside it. Additionally, participants could delete translations from the history by clicking the cross button, though we found that participants rarely used this feature in our study.
3.1.3. Edit Profile Dialog Box
The edit profile dialog box was accessible by clicking a button on the page header (Figure 1E). This displayed a short form (Figure 1[II]) containing information about the participant including their background, age, hobbies, location, and favorite food. This profile information was used to tailor the translation to the participant’s context. The study consisted of two sessions: in the first, participants could not see or edit their profiles, while in the second, they were allowed to update their profiles to further tailor the personalization.
@@ 3.1.4. Finishing a Session
Once participants were satisfied with their understanding of the scientific material, they were instructed to click the finish button (Figure 1F) in the top right corner. This action triggered a pop-up (Figure 1[III]) that displayed a thank you message and an additional question asking if they would like to receive the original research paper PDF via email after the study. This served as a proxy measure for their interest in the content after reading the personalized translations.
3.2. Implementation and Prompt Design
We built the interface using React, a popular JavaScript library for front-end development, and implemented the backend in JavaScript. When a participant requested a translation at a specific personalization degree, an API call was made to OpenAI’s GPT-4o model to generate the translation. As shown in Figure 2, we used a detailed prompt template to guide the model’s output.
Figure 2. Prompt design for generating personalized translations. The prompts (A, B, C, and D) are combined in order, and then sent to LLM through an API call to generate tailored translations of scientific articles.
3.2.1. LLM Prompt
The prompt began with a high-level task overview, instructing the model to use analogies to explain a science article to general audiences. It then defined the personalization spectrum, providing clear descriptions for degrees 0, 50, and 100. Following this, the prompt included in-context examples of both generic and personalized translations. To improve quality, we applied chain-of-thought prompting, guiding the model to reason through its responses. Finally, the model was given the scientific article and asked to produce a two-paragraph translation while avoiding titles, honorifics, or other formalities. The exact prompt template is given in Figure 2.
We arrived at this prompt template through iterative experimentation. Once the research team was satisfied with the LLM’s output, we conducted a pilot study with three research scientists who had published their work (in Computer Science, Biomedical Engineering, and Electrical Engineering, respectively) in the past three years to gather feedback on personalization quality and to confirm that the model was not hallucinating content in its translations. We used their own research papers in this pilot study to let them verify the personalized translations. We proceeded with the main study after receiving this confirmation from all three of them.
@@ 4. User Study
Our study aimed to obtain readers’ feedback and usage data regarding two aspects of the tool: i) the utility of AI-personalized translation of scientific text (RQ1), and ii) the impact of interactive reading features on user experience (RQ2). The study consisted of two sessions, designed to gradually introduce tool features to participants. In the first session, participants interacted with TranSlider that implicitly personalizes scientific articles based on their user profiles, similar to how web browsers and social media platforms display personalized content based on collected user metadata (e.g., Google Chrome’s personalization through Google accounts). The second session introduced a more full-featured interface where participants have the full autonomy to explicitly edit their profiles to manipulate the direction of personalization. Throughout both sessions, participants control a slider to explore different degrees of personalization along a continuous spectrum. We describe the detailed study procedure in Section 4.3.
@@ A P1 Female 43 UX Design 3 Yes — No health environment
We recruited 15 participants (10 female and five male) from our organization through internal mailing lists and word-of-mouth referrals. While our organization employs highly educated individuals with an analytical mindset, we used a diverse recruitment strategy to broaden the applicability of our study. Specifically, to ensure that our sample represented general readers rather than scientists or science writers, we excluded individuals with a doctoral degree or equivalent job titles (e.g., research scientists). All participants were fluent in English and aged 18 or above (range: 23-43, M=26.7, SD=5.3).
Participants came from diverse educational backgrounds, including Business, Economics, Industrial Engineering, Computer Science, Information Systems, Cognitive Science, Biology, and Statistics. We asked about their frequency of reading scientific articles (e.g., science blogs) and experience with LLMs (e.g., ChatGPT). Nine participants reported actively reading scientific texts at least once a month, while three had not engaged with any scientific articles in the past three years. All participants had prior experience with LLMs, with eight actively using them to comprehend complex texts like technical reports or academic papers. However, four participants were familiar with LLMs but not actively utilizing them daily. To further characterize our sample, we collected information on participants’ highest level of education. Table 1 provides detailed demographic data.
@@ 4.3. Procedure
The study was conducted 1-on-1 and in-person. We audio and screen recorded study sessions that took an hour on average (M: 60.5 min, SD: 10.9 min). All participants received a 75 USD virtual gift card as compensation for their time.
@@ 4.3.4. Debrief Interview (20 min)
After completing both sessions (implicit and explicit), we conducted a semi-structured debrief interview aligned with our research questions. The questions included i) the participants’ perceived quality and utility of AI-personalized translations (RQ1), ii) their experience with interactive reading features such as multiple translations, history items, and the slider (RQ2), iii) any particular aspects they liked or disliked, and iv) overall concerns and suggestions for improvement. At the end, participants were free to ask any questions or share additional comments.
@@ 4.4.1. User Behavioral Logs
We recorded user logs to analyze participants’ tool exploration behaviors. The collected data includes user profile information (implicit), profile edits (explicit), explored degrees of personalization, all generated translation texts, favorite translations marked by users, deleted translations, and precise timestamps for all these actions. Then we analyzed them by plotting and computing descriptive statistics (e.g., mean, std, max, and min). We did not conduct a statistical analysis of our data between conditions or groups, because we had a small sample size ($N=$15) (Gelman and Carlin, 2014). More importantly, our study design was intended to enable qualitative analysis.
@@ 4.4.2. Interview Logs
All interview recordings were transcribed using automated transcription services. We conducted thematic analysis (Braun and Clarke, 2006) over the transcribed quotes that were organized by interview questions for each participant. For the open-coding process, the first author read quotes and labeled them with initial codes (i.e., low-level codes) using Miro board cards [^2] https://miro.com/. Then the first and the second authors collaboratively reviewed the low-level codes to discuss and group codes into potential themes (i.e., high-level themes). The first author structured a thematic map showing themes for each corresponding research question. Next, the research team collaborated to refine the thematic map more clearly, build a coherent narrative, and reach a consensus. During this process, we excluded themes that were not directly relevant to our research questions. Ultimately, we finalized the themes for each research question: the utility of AI-personalized translations (RQ1), the impact of interactive reading features on user experience (RQ2), and proposed use cases and associated concerns.
@@ 5. Findings
We first set the context by summarizing user behavior using log data from TranSlider. Then we synthesize participants’ interview responses supported by user log data to substantiate additional insights and suggest potential opportunities for future research directions. Our findings cover three key aspects: i) the utility of AI-personalized translation for scientific text, ii) the impact of interactive reading features on user experiences, and iii) the potential use cases and concerns.
@@ 5.1.1. Users’ Exploration Behaviors
Participants generated a total of 268 AI-personalized translations using TranSlider, with an average of 8.93 translations explored per session (min: 5, max: 19). We observed a positive correlation (Pearson’s $r=0.36$) between the degree of personalization and length of the translation (see Figure 3), showing that more personalized translations tended to be slightly longer. This subtle increase in length could be attributed to the need for integrating more detailed personal context in highly personalized translations, whereas low personalization required little to no incorporation of personal information.
Participants demonstrated various exploration trajectories, including: i) incremental—starting from low degrees like 2 and moving toward higher degrees, ii) decremental—starting from high degrees like 98 or 100 and moving toward lower degrees, and iii) edge degrees—exploring the extremes (0 or 100) before exploring moderate degrees. Regardless of their exploration starting points, participants often revisited and compared previous translations, moving back and forth between different personalization degrees. While participants largely explored edge degrees more (0 and 100), many of them eventually favored translations with moderate degrees (see 4(a)). The mean value of participants’ favorite translations was 53.14 (SD: 33.24, Median: 51.0), indicating a tendency towards moderate personalization rather than extremes such as hyper-personalization or none at all.
@@ 5.1.2. Motivations for Reading Scientific Articles: Extrinsic and Intrinsic
We categorize participants’ motivation for scientific articles as extrinsic and intrinsic based on their pre-study interview responses. Those with extrinsic motivations (N=7) read scientific articles primarily when required for work-related purposes, while intrinsically motivated individuals (N=8) read them voluntarily, even in their spare time and regardless of work relevance. 4(b) illustrates the distribution of participants’ preferred translation degrees within each motivation category. Descriptively, we see that participants with extrinsic motivation tend to prefer a lower degree of personalization (M: 40.52, SD: 29.84), while those with intrinsic motivation favor higher degrees (M: 62.75, SD: 29.35). While our study was not designed to compare statistical differences between these two groups (Montgomery et al., 2018), this descriptive result shows that motivation could be an important factor in designing personalization interfaces.
(a) Participants tend to explore extreme personalization degrees (0 or 100) more frequently, yet their preferred translations typically fell within moderate ranges, despite less exploration of middle ones.
(b) Those with extrinsic motivation for reading scientific articles tend to prefer personalization lower degree, while those with intrinsic motivation liked translations with higher degrees.
Figure 4. (a) shows the histogram of two distributions: all translations vs. user-selected favorite translations. (b) displays a box plot of favorite translations by different motivations for reading scientific articles.
@@ 5.2. Utility of AI-Personalized Translation for Scientific Text (RQ1)
All participants reported the benefits of personalized translations for scientific articles. However, individuals preferred varying degrees of personalization for different reasons. We asked participants to explain their preferences and rationale with concrete examples of translations they had explored. Based on these examples, we categorized their responses into two groups: those preferring higher degrees of personalization (ranging from 51 to 100) and those preferring lower degrees (ranging from 0 to 45).
@@ 5.2.1. Participants Who Preferred Higher Degree Liked Relatable and Informative Translations
Participants who preferred translations with higher degrees of personalization appreciated that these highly relatable explanations facilitated their understanding of complex ideas (e.g., P7, P8, P9, P10, P14). They further stated that these personalized translations, which extensively incorporated participants’ personal contexts, resonated with them and were funny. For instance, P10 appreciated the translation began with ‘running’, one of her hobbies, making challenging scientific concepts more relatable and easier to understand. Similarly, P14 found some personalizations funny as they incorporated the safety and security aspects of his hometown and his cultural background.
@@
Figure 5. Several examples of preferred translations. While each personalized translation is longer (two paragraphs), we present specific parts that participants pointed out to highlight why they liked or disliked the passage.
@@
P14: I feel like this last point was pretty good. Because it says, I’m not sure if the other ones really talked about it that much, but that the cleaner energy could be causing pollution. So the implication of the study is kind of good [useful].

### v2

@@ 4.1.1. Translation Panel
The translation panel has the main personalization features (Figure 1 A , B , C ). The personalized title of the scientific article being viewed is displayed at the top with the personalized translation underneath it. A slider, labeled from weak to strong personalization, allowed participants to adjust the degree of personalization. The current personalization degree, a numeric value between 0 and 100, is shown just beneath the slider. Participants were instructed to adjust the slider to their preferred degree value before translating using the Translate button. This action triggers an LLM call, which generates and displays a personalized translation.
@@ 4.1.2. History Box
The history box on the right allows participants to revisit, review, and compare translations they had seen during the session. Each new translation is automatically added to the stack, labeled with its personalization degree and the time of generation. Users can click on any past translation to view it, which would update the translation panel accordingly. They could also mark a translation as a favorite by clicking on the star icon beside it. Additionally, users can delete translations from the history by clicking the cross button, though we found that participants rarely used this feature in our study.
4.1.3. Edit Profile
The edit profile dialog box is accessible by clicking a button on the tool header (Figure 1  E ). This displays a short form (Figure 1 [II]) containing information about the participant including their background, age, hobbies, location, and favorite food. This profile information is used to tailor the translation to the user’s context.
@@ 4.1.4. Finishing a Session
The finish session button (Figure 1  F ) served as a proxy measure for their interest in the content after reading the personalized translations. The button triggers a pop-up (Figure 1 [III]) that collects user response on whether they would like to receive the original research paper PDF via email. This button was used at the end of each article’s translation session.
4.2. System Workflow and Implementation
The workflow of TranSlider is illustrated in Figure 2. Once a user adjusts the personalization slider and submits a translation request on the frontend, the backend integrates both the selected personalization degree and the original scientific text into the prompt template (detailed in Figure 3). Inspired by existing LLM-powered tools’ prompt designs (e.g.,  (Chung and Kreminski, 2024)), we used the chain-of-thought style instruction (Wei et al., 2022) to guide the model to reason how the slider input value (personalization degree) should be reflected in the output (personalized translation). With this combined prompt input, the backend makes an API call to the LLM to generate the personalized translation. The API response with the AI-personalized translation is then transmitted back to the translation panel in the frontend. We implemented TranSlider using React.js for the frontend interface and Node.js for the backend. We used OpenAI’s GPT-4o model for the LLM.
Note on LLM Model Choice
While our implementation utilized GPT-4o, this model could be readily replaced with alternative commercial or open-source models. For instance, specialized models like OpenScholar-8B—specifically designed for scientific contexts through retrieval augmentation—outperform GPT-4o for scientific content (Asai et al., 2024) and could potentially generate higher-quality personalized translations. Conversely, substituting with older or less capable models such as GPT-Neo (Black et al., 2021) would likely result in lower quality and unreliable translations. Our preliminary testing confirms that alternative models like Meta’s Llama-3 (open source) (Meta, 2024), Anthropic’s Claude 3.5 (commercial) (Anthropic, 2024) also generate comparable quality of translations, suggesting that our tool could incorporate different model architectures. Whether there would be any noticeable distinctions in the overall translation quality depending on the model choice is left for future investigation.
Figure 2. The workflow of TranSlider.
Figure 3. Prompt Template Structure. The prompt begins with a Task overview, instructing the model to use analogies to explain a science article to general audiences. The prompt context consists of four parts: 1. The user profile; 2. Personalization spectrum with descriptions for degrees 0, 50, and 100; 3. In-context examples of both generic and personalized translations; 4. Original scientific article content with the slider value.
4.3. Pilot Study
Once the research team was satisfied with the LLM’s output, we conducted a pilot study to determine whether the translations were understandable and free from hallucinations or misinformation. We recruited three research scientists outside of our research team. They all had published their work in the past three years. We used their authored publications in Computer Science, Biomedical Engineering, and Electrical Engineering, respectively, for the verification process.
4.3.1. Procedure
The first author conducted this study in a one-on-one in-person setting. Upon participants’ arrival, we introduced TranSlider as a tool designed to help general audiences understand complex scientific texts. We explicitly stated that the purpose of this pilot study was to verify whether the AI-personalized translations were comprehensible and did not hallucinate any misinformation. We then asked participants to provide a link to their recent publication that they would like to share with the general audience. After a brief tutorial of our tool, we had them explore and interact with TranSlider at their own pace without any time constraints. We proceeded to the debriefing interview — the overall impression of the personalized translations and whether they had seen any misinformation or hallucinations — only when participants indicated they were ready to share their thoughts. Each of the three pilot studies took about 30 minutes to complete.
4.3.2. Results
All three scientists found the translations surprisingly understandable. They also confirmed that they could not detect any hallucinations or misinformation in the translations they reviewed. The electrical engineering researcher provided additional feedback on two aspects: the topic selection and the structure of translations. First, they noted that a scientific article’s subject matter could influence its perceived benefits to general audiences. Some topics, like health and nutrition, might naturally interest the general audience, while they doubted whether some other topics, like environmental or geological sciences, would draw similar interest. Secondly, they suggested that highlighting some practical implications of studies could make translations more directly beneficial to general audiences.
4.3.3. Implications
The results of this pilot study influenced our prompt and study design. We refined the instruction prompt (see 4. in Figure 3) to generate two distinct paragraphs: the first presenting an analogy explaining the study, and the second discussing implications to the user (example in  Figure 4). Previously, the translation only explained the study without offering personalized implications. We also decided to include two different types of scientific articles for our study: one with directly relatable content and another with less relatable content (Section 5.1).
Figure 4. An example translation (degree = 100) of P14, whose hobby is basketball and rock climbing, and lives in San Francisco, CA. Their favorite foods are Korean barbeque and beef noodle soup. The first paragraph gives a brief explanation of the scientific content based on their personal context: location (San Francisco) and hobby (Basketball). The second paragraph presents the implications of this scientific knowledge to the reader’s diet and health, saying that both Korean BBQ and beef noodle soup have a lot of sugar.
@@ 5. User Study
Our study aimed to obtain readers’ feedback and usage data regarding two aspects of the tool: i) the utility of AI-personalized translation of scientific text (RQ1), and ii) the impact of interactive reading features on user experience (RQ2). Below, we describe the details of our study design, including the choice of science articles, participants, study procedure, and data analysis methods.
@@ B P8 Male 25 Economics 3 Yes — No environment health
We recruited 15 participants (10 female and five male) through mailing lists and word-of-mouth referrals. Our recruitment criteria were designed to represent the general public with an interest in science, but not scientists. We recruited people above 18 years of age and different educational domains—a mix of science-based (e.g., engineering) and non-science based fields (e.g., business). We excluded people who: (a) had a doctorate degree or a research position, and (b) had a degree in health and environment fields—the topics of the scientific articles used in our study. This approach allowed us to secure participants with various domain backgrounds with a baseline level of interest in science and unfamiliarity with the chosen articles’ fields. We also asked participants how often they read scientific articles (e.g.,  science blogs) to ensure representation across a spectrum of regular readers and those who rarely engaged with scientific content.
Table 1 illustrates the details of participants’ demographics. All participants were fluent in English and aged between 23-43 years (M=26.7, SD=5.3). Most participants held bachelor’s degrees in various fields, while one participant (P1) had no degree. In the past three years, eight participants read scientific articles (e.g.,  science blogs) at least monthly, while the others did so rarely. Regarding LLM experience, all participants were familiar with tools like ChatGPT; four actively used them to understand complex texts (e.g.,  technical reports or academic papers), while the others did not.
@@ 5.3. Procedure
The study was conducted one-on-one and in-person. We audio and screen recorded study sessions that took an hour on average (M=60.5 min, SD=10.9 min). All participants received a 75 USD virtual gift card as compensation for their time.
@@ 5.3.4. Debrief Interview (20 min)
After completing both sessions (implicit and explicit), we conducted a semi-structured interview aligned with our research questions. The questions included i) the participants’ perceived quality and utility of AI-personalized translations (RQ1), ii) their experience with interactive reading features such as multiple translations, history items, and the slider (RQ2), iii) any particular aspects they liked or disliked, and iv) overall concerns and suggestions for improvement. At the end, participants were free to ask any questions or share additional comments.
@@ 5.4.1. User Behavioral Logs
We recorded user logs to analyze participants’ tool exploration behaviors. The collected data includes user profile information (implicit), profile edits (explicit), explored degrees of personalization, all generated translation texts, favorite translations marked by users, deleted translations, and precise timestamps for all these actions. We analyzed them by plotting and computing descriptive statistics (e.g., mean, std, max, and min). We did not conduct a statistical analysis of our data between conditions or groups, because we had a small sample size (N=15) (Gelman and Carlin, 2014). More importantly, our study design was intended to enable qualitative analysis.
@@ 5.4.2. Interview Logs
We used automated services to transcribe all interview recordings. Then we conducted thematic analysis (Braun and Clarke, 2006) on the quotes organized by interview question for each participant. For the open-coding process, the first author read quotes and labeled them with initial codes (i.e., low-level codes) using Miro board cards [^3] http://miro.com/. Then the first and the second authors collaboratively reviewed the low-level codes to discuss and group codes into potential themes (i.e., high-level themes). The first author structured a thematic map showing themes for each corresponding research question. Next, the research team collaborated to refine the thematic map more clearly, build a coherent narrative, and reach a consensus. During this process, we excluded themes that were not directly relevant to our research questions. Ultimately, we finalized the themes for each research question: the utility of AI-personalized translations (RQ1), the impact of interactive reading features on user experience (RQ2), and proposed use cases and associated concerns.
@@ 6. Findings
We first summarize findings from user behavior log data. We then synthesize participants’ interview responses supported by user log data to substantiate additional insights and suggest potential opportunities for future research directions. Our findings cover three key aspects: i) the utility of AI-personalized translation for scientific text, ii) the impact of interactive reading features on user experiences, and iii) the potential use cases and concerns.
@@ 6.1.1. Users’ Exploration Behaviors
Participants generated a total of 268 AI-personalized translations using TranSlider, with an average of 8.93 translations explored per session (min: 5, max: 19). We observed a positive correlation (Pearson’s $r=0.36$) between the degree of personalization and length of the translation (Figure 5), showing that more personalized translations tended to be slightly longer. This subtle increase in length could be attributed to the need for integrating more detailed personal context in highly personalized translations, whereas low personalization required little to no incorporation of personal information.
Participants demonstrated various exploration trajectories, including: i) incremental—starting from low degrees like 2 and moving toward higher degrees, ii) decremental—starting from high degrees like 98 or 100 and moving toward lower degrees, and iii) edge degrees—exploring the extremes (0 or 100) before exploring moderate degrees. Regardless of their starting points, participants often revisited and compared previous translations, moving back and forth between different personalization degrees. While participants largely explored edge degrees more (0 and 100), many of them eventually favored translations with moderate degrees (6(a)). The mean value of participants’ favorite translations was 53.14 (SD=33.24, Median=51.0), indicating an average tendency towards moderate personalization rather than extremes such as hyper-personalization or none at all. However, the high SD shows that participants prefer translations across a wide range of degrees, rather than clustering around a central tendency. This supports our original motivation that singular translations of science articles may not be enough to capture diverse audiences.
(a) Participants tend to explore extreme personalization degrees (0 or 100) more often, yet their preferred translations typically fell within moderate ranges, despite less exploration of middle ones.
(b) Those with extrinsic motivation for reading scientific articles tend to prefer personalization lower degree, while those with intrinsic motivation like translations with higher degrees.
Figure 6. Illustration (a) shows the histogram of two distributions: all translations vs. user-selected translations. Illustration (b) displays a box plot of favorite translations by motivations for reading scientific articles.
@@ 6.1.2. Motivations for Reading Scientific Articles: Extrinsic and Intrinsic
We categorize participants’ motivation for scientific articles as extrinsic and intrinsic based on their pre-study interview responses. Those with extrinsic motivations (N=7) read scientific articles primarily when required for work-related purposes, while intrinsically motivated individuals (N=8) read them voluntarily, even in their spare time and regardless of work relevance. 6(b) illustrates the distribution of participants’ preferred translation degrees within each motivation category. Descriptively, we see that participants with extrinsic motivation tend to prefer a lower degree of personalization (Mean=40.52, SD=29.84), while those with intrinsic motivation favor higher degrees (Mean=62.75, SD=29.35). While our study was not designed to compare statistical differences between these two groups (Montgomery et al., 2018), this descriptive result [^4] We emphasize that our quantitative analysis is limited by the small sample size and participant homogeneity regarding education levels. Hence, it should be treated as preliminary. We encourage a robust analysis before such a system is deployed. shows that motivation could be an important factor in designing personalization interfaces.
@@ 6.2. Utility of AI-Personalized Translation for Scientific Text (RQ1)
All participants reported the benefits of personalized translations for scientific articles. However, individuals preferred varying degrees of personalization for different reasons. We asked participants to explain their preferences and rationale with concrete examples they had explored. Based on these examples, we categorized their responses into two groups: those preferring higher degrees of personalization (ranging from 51 to 100) and those preferring lower degrees (ranging from 0 to 45).
@@ 6.2.1. Participants Who Preferred Higher Degree Liked Relatable and Informative Translations
Participants who preferred translations with higher degrees of personalization appreciated that these highly relatable explanations facilitated their understanding of complex ideas (e.g., P7, P8, P9, P10, P14). They further stated that these personalized translations, which extensively incorporated participants’ personal contexts, resonated with them and were funny. For instance, P10 appreciated the translation began with ‘running’, one of their hobbies, making challenging scientific concepts more relatable and easier to understand. Similarly, P14 found some personalizations funny as they incorporated the safety and security aspects of their hometown and cultural background (see translation of P14 in Figure 7).
@@
Figure 7. Several examples of preferred translations. While each personalized translation is longer (two paragraphs), we present specific parts that participants pointed out to highlight what they liked or disliked.
@@
P14: I feel like this last point [implication] was pretty good. Because it says, I’m not sure if the other ones really talked about it that much, but that the cleaner energy could be causing pollution. So the implication of the study is kind of good [useful].